OpenStudy (anonymous):

A car can travel 25 miles per gallon. Approximately how many gallons of fuel will the car need to travel 30 km?

OpenStudy (abhisar):

Hi! What do you think it should be? Have you tried solving it urself first?

OpenStudy (anonymous):

i dont know how

OpenStudy (abhisar):

ok, the car can do 25 miles in one gallon of fuel then how many gallon it needs to do just one mile?

OpenStudy (anonymous):

1.6 km

OpenStudy (abhisar):

Ummm.... No I am asking how much gallons it need to do one mile. 25 miles need 1 gallon so 1 mile will need 1/25 = ? gallon?

OpenStudy (anonymous):

idk this

OpenStudy (aaronandyson):

Hint : 1 mile = 1.6 km Therefore, 25 time 1.6 = 40

OpenStudy (aaronandyson):

Therefore, 40 times x = 30 Therefore, 40x = 30 Therefore,x = 30/40 Therefore,x = 3/4

OpenStudy (aaronandyson):

Therefore, x = 0.75 :)!

OpenStudy (aaronandyson):

Go through my calculation and see if you haven't understand anything..

OpenStudy (anonymous):

AaronAndyson, are you a college instructor?

OpenStudy (anonymous):

Nope you are not lol You have the name of a college instructor I know :-)

OpenStudy (anonymous):

thank uuuu

OpenStudy (aaronandyson):

:) @Nixy